OpenStudy (anonymous):

If a number is decreased by its square, the result is 2/9. Find the number

OpenStudy (shane_b):

This is the same type of problem sakigirl, just write the equation out and solve it similarly:\[x-x^2=\frac{2}{9}\]

OpenStudy (anonymous):

let x be that number a number decreased by its square is x - x^2 = 2/9 9x - 9x^2 = 2 9x^2 - 9x + 2 = 0 (3x-2)(3x-1) = 0 equate each factor to 0 then solve for x

OpenStudy (anonymous):

I get -1 and +2

OpenStudy (anonymous):

remember you're looking for x. you'll get 3x = 2 and 3x = 1. You have to divide by the coefficient to get x

OpenStudy (anonymous):

Wait, this is what I did: x-x^2-2/9=0 I multiplied everything by nine to make it easier on me: 9x^2-9x-18 And then I factored out the 9: 9(x^2-x-2) And then I factored from there on: (x-2)(x+1) which is how I get my answer: x=2, -1

OpenStudy (anonymous):

if you multiply 2/9 by 9 you get 2. 9 just cancels out

OpenStudy (anonymous):

not 18

OpenStudy (anonymous):

Oh gosh, what a careless mistake. So I guess I have to use the quadratic formula?

OpenStudy (anonymous):

you can. but factorable too. i presented the factors,

OpenStudy (anonymous):

but that kind of factoring is confusing so quadratic formula may be better

OpenStudy (anonymous):

Okay thank you!